# Fuzzy Ontology Embeddings and Visual Query Building for Ontology Exploration

### v2

@@ Abstract
The visual interface supports browsing, query composition, and partial search without requiring formal syntax. By combining fuzzy semantics with embedding-based reasoning, FuzzyVis enables flexible interpretation, efficient computation, and exploratory learning. Case studies demonstrate how FuzzyVis supports subtle information needs and helps users uncover relevant concepts in large, complex ontologies.
@@ 1 Introduction
To implement our querying approach, we represent ontologies as fuzzy ontologies with fuzzy interpretations, where each concept is treated as a fuzzy set assigning a membership degree to each element in the interpretation domain. Given an ontology, we use existing fuzzy reasoners—such as fuzzyDL [9], DeLorean [48], and FALCON [58]—to construct such interpretations that approximately satisfy the ontology’s axioms. Based on these interpretations, we introduce a novel method for generating ontology embeddings, where each concept is represented as a vector of membership degrees over a fixed, ordered set of domain elements. This allows us to precompute embeddings for all primitive concepts offline. At query time, when a user defines a composite concept using conjunction, disjunction, or negation over known primitives, its embedding is computed by applying the corresponding fuzzy operations element-wise to the membership vectors. This compositional capability enables flexible query construction and is not supported by existing ontology embedding methods. The resulting embeddings also support efficient similarity computations (e.g., cosine similarity) between user-defined concepts and existing primitive concepts in the ontology, enabling approximate matching even when no exact concept exists.
@@ 3.1 Ontologies: Syntax and Semantics
In our setting, we are more concerned with querying ontologies to retrieve individuals or concepts that match a user’s information needs. Common query types include conjunctive queries (CQs), unions of conjunctive queries (UCQs), and recursive extensions that support reachability-like conditions. DL query languages such as SPARQL-DL [54], nRQL [30], and QueryTab [49] combine logical reasoning with structured query syntax. However, these languages require familiarity with formal syntax and detailed knowledge of the ontology’s structure, and their evaluation is often computationally expensive. This limits their suitability for users seeking intuitive, flexible, and interactive exploration.
@@ 4 Ontology Querying using Fuzzy Ontology Embeddings
Scalable and optimized retrieval: Since all concept representations are vectors, similarity search over primitive concepts can be delegated to optimized vector databases (e.g., FAISS [37], Qdrant, or Chroma). These systems support fast top- retrieval and efficient indexing, enabling responsive and semantics-aware retrieval for exploratory query interfaces.
4.1 Constructing Interpretations for Hierarchical Ontologies
In many practical domains such as biomedicine or botany, ontologies take the form of hierarchies or trees, where each concept has parent concepts, some number of children, and there are no cycles (e.g., HPO [29], PO [36]). For such ontologies, we can construct a synthetic but semantically meaningful fuzzy interpretation that captures hierarchical relationships and allows us to generate expressive embeddings efficiently.
@@ 4.1 Constructing Interpretations for Hierarchical Ontologies
This method provides an effective and interpretable way to construct embeddings from hierarchical ontologies, capturing both local concept similarity and the global structure. It also allows control over smoothness and generalization via the decay parameter , making it a practical tool for building fuzzy ontology embeddings in common tree-structured settings. We refer to these synthetic embeddings as -embeddings.
@@ 5.1 FuzzyVis Architecture and Implementation
The system workflow begins with the user uploading ontology data in standard formats such as .owl or .ttl. The ontology is then processed and stored on the back-end. Once loaded, users can explore the ontology through interactive views and visualizations on the front-end. This exploration is supported by panels and controls that allow users to manipulate views, adjust parameters, and construct queries.
As users interact with the system, in the front-end FuzzyVis dynamically loads relevant ontology data and discards unused portions to optimize memory usage and ensure scalability for large ontologies. Many ontologies are so large that users cannot feasible view them in their entirety. By discarding unused portions users’ systems do not waste unnecessary resources. This optimization ensures that exploration of the ontology remains responsive. To maintain this responsiveness during querying, the back-end handles query resolution and returns the results to the query building panels of the front-end. Users can explore these results to navigate the ontology or refine subsequent queries.
While the front-end handles user interaction and visualization, all computations—including query evaluation and similarity search—are performed on the back-end. Embeddings for primitive concepts are stored in a Chroma vector database, 111https://github.com/chroma-core/chroma as described in Section 4. This enables quick query resolution, especially for ontologies with thousands of concepts, and or large embeddings.
@@ 5.2 Back-End Server
The back-end consists of four main components: the ontology loader, ontology-database, vector-database, query resolver. The ontology loader handles uploaded ontologies—pre-uploaded or submitted via the front-end. It parses ontology files and stores their contents in an ontology-database using the OwlReady2 python library [40]. During this process, additional metadata and statistics—such as subtree sizes and number of children—are computed and stored. The vector-database stores the embeddings for each primitive concept. The vector-database is separate from the ontology-database to perform quick similarity calculation and top-k retrieval. Our implementation uses Chroma for its vector-database. Just like for ontologies, users can submit embedding vectors or use pre-uploaded ones. Uploaded embedding vectors can be generated using whatever means users desire—such as fuzzy ontology reasoners—as long as the vectors have the properties of those we define in section 4. If users do not provide an embedding the system will generate a simple embedding using leaf concept membership and subsumption. The selection and loading of embeddings is also handled by the ontology loader. Lastly, the query resolver computes embeddings for user-defined composite concepts. The vector-database uses these composite concept embeddings to retrieve similar primitive concepts. A FLASK server API handles Communication between the back- and front-end. The front-end requests ontology information and the back-end provides.
@@ 5.3 Front-End Interface
The front-end is composed of several regions consisting of panels and controls that support interactive ontology exploration. Figure 2 shows an overview of the front-end interface as seen by the user. The primary visualization takes up the central region of the interface and is the users’ primary means of navigating the ontology (Figure 2.A). Users can change between different visualizations—treemps, nested lists, and network graphs—in this central region. The second region (Figure 2.B), the top header, provides quick access for concept search, switching between loaded ontologies, and loading ontologies. Users can manipulate the behavior and control the presentation using the visualization controls in the leftmost region (Figure 2.C). The first of the rightmost regions (Figure 2.D) is the concept panels where users can store concepts for future use and look at concept details. The second rightmost region (Figure 2.E) is the query building panels where users define queries and examine their results. FuzzyVis ’s front-end is compatible with all major web browsers.
@@ 5.3.1 Primary Visualizations
FuzzyVis ’s central region is contains the primary visualization (Figure 2.A). Here users see a representation of the ontology to aid them in exploration and understanding. Each component of FuzzyVis serves to modify, control, and navigate the primary visualization, with the back-end providing any information needed by the primary visualization. The primary visualization can take the form of any suitable representation of the ontology. Visualization should meet the following criteria:
@@
Classical visualizations such as nested lists, icicle charts, network graphs, and treemaps can only partially meet these requirements. In fact, no single static visualization can satisfy every requirement. Network graphs often become unreadable with high volumes of concepts, nested lists poorly use space, and even treemaps struggle to convey secondary features. To address this problem, part of our solution is to supplement the primary visualization with secondary panels that provide information that the primary visualization cannot. The other part of the solution is to utilize dynamic interactive visualizations. To illustrate how we meet these requirements and explain the functionality of FuzzyVis we will discuss the nested treemap version of the primary visualization in this section.
Nested treemaps consist of nested rectangles, where each rectangle corresponds to a concept. Parent-child relationships are represented by the child concept’s rectangle being positioned inside the parent concept’s area. Nested treemaps are a suitable visualization of ontologies as they meet the majority of the requirements. In particular, such visualizations clearly distinguish concepts, are consistent, space-efficient, and structure-revealing. By extending nested treemaps to be dynamic and interactive, secondary properties can be shown and egregious concept density can be mitigated. An image of FuzzyVis ’s nested treemap, henceforth referred to as the treemap, is shown in Figure 3. As showing the entirety of ontologies is impractical, we implemented the treemap to only show concepts within an adjustable distance of a selected concept. As users interact with the treemap it adjusts to show the selected concepts, allowing users to navigate the ontology. Furthermore, the treemap has a number of additional features, interactions, and integration with the rest of FuzzyVis ’s front-end that aid users in their exploration activities.
@@
Selecting concepts in the treemap refocuses the visualization onto them. By selecting concepts, users drill down into concept’s subtrees or up to through the parent—navigating the ontology. Beyond selection, users are able to drag concepts to other components, and pin concepts to mark them. Supplementary visual marks placed by users help them in future navigation by acting as signpost to aid in orientation. Furthermore, in cases where a concept is present in the subtrees of disjoint parents a visual mark present in multiple locations helps users identify this. For example, in HPO the Cardiac valve calcification concept is both a child in the Cardiovascular calcification and Abnormal heart morphology subtrees. When pinned Cardiac valve calcification shows up in both areas indicating to users overlap between the parent concepts. Without visual aids users need to manually read each concept label in order to identify them. Afterwards, once users have collected concepts that interest them, they can create queries using the Query Building Panels to direct their future exploration.
Users exploring ontologies may want to focus on a specific concept and see its content as though it were the ontology’s root. Simply increasing the depth quickly leads to two issues. One, child concepts a few layers deep have a small amount of space available to them; users must mouse over these children to see a tooltip to identify them. Two, when every subtree is full of child concepts the visualization becomes very dense, overwhelming users.
@@
To address these issues, we implemented a toggle for a focus mode that enables on-the-fly visual distortions. During focus mode, when users mouse over concepts, FuzzyVis applies a discrete Cartesian distortion (often referred to as a fisheye distortion (or lens)), greatly increasing the moused-over concept’s area while maintaining its relative position to the rest of its siblings. Within the focused concept’s increased area, the subtree extends to show the children of the focused concept. Users that mouse over the added children can continuously drill down until there is insufficient space for further expansion. If users want to explore multiple subtrees simultaneously they can lock a concept as a locus and the distortion will maintain multiple focal points. Moreover, focus mode is not limited to extending the treemap deeper, as any focused locus can be transformed into a different visualization. For example, a concept cell can be converted to a node network graph or a panel showing class details.
@@
As an example of the focus mode feature, in Figure 3, the depth is set to two and the Abnormal nervous system physiology is selected. Normally, no concepts other than the children of Abnormal nervous system physiology are visible under these conditions. Yet, with focus mode turned on, the user mousing over Abnormal central motor function fills the subtree with child concepts. Afterwards, when the user mouses over the child Abnormality of coordination further expands the subtree to show this concept’s children (e.g, Slurred speech). This could have continued until the user reached a leaf concept or ran out of screen space, but, the user decided this depth was sufficient and locked Abnormality of coordination as a locus. Thus, focus mode provides users with a dynamic and space-efficient way to quickly explore deeper into an ontology while maintaining context and their position. Furthermore, dynamic distortions are applicable to other kinds of visualizations, such as network graphs and charts.
Users control the presentation and behavior of the primary visualization through the Visualization Controls (discussed in detail in section 5.3.3). For the treemap, users can adjust the displayed depth, extend the labels with more information (e.g., depth numbers), control how concepts are tiled, which concepts are shown, and highlight sections of the visualization using these settings. Giving users control over the visualization options is necessary, as some ontologies present better under different conditions. For example, in HPO, among the children of Phenotypic abnormality there is Abnormality of the musculoskeletal system and Abnormality of the voice. Abnormality of the voice contains few classes and is small and almost unnoticeable in the treemap when cell scaling is set to be proportional to subtree size and no normalization is applied. Abnormality of the musculoskeletal system, on the other hand, takes up a significant amount of space. Users looking for conditions related to the voice may assume that HPO lacks them if they explored the ontology with the prior settings. With the cell scaling set to equal, each child concept of Phenotypic abnormality takes up the same amount of space. Users see both Abnormality of the musculoskeletal system and Abnormality of the voice and are less likely to miss important concepts. Furthermore, users can toggle these options to see how the visualization changes—toggling the equal scaling on and off shows users which concepts take up the majority of an ontology while quickly returning to a view suitable for navigation. The ability to adjust the visualization lets users see more of an ontology’s information without restricting them to suboptimal visualizations.
@@
FuzzyVis’s primary visualization helps users to explore the structure and semantics of ontologies. In conjunction with other front-end panels, it provides a high-level spatial overview that enables users to identify key concepts and regions of interest. By presenting concepts with a dynamic visualization FuzzyVis supports rapid navigation aiding in users mental model formation. This visual aid is necessary when working with complex knowledge domains where manual inspection of concepts would be inefficient or overwhelming. Ultimately, the primary visualization serves as the main interface through which users interact and explore the ontology.
@@ 5.3.2 Header Bar
FuzzyVis’s header bar (Figure 2.B) has three interactive elements. Users can load ontologies, switch between loaded ontologies, and perform keyword search to quickly navigate using these interactive elements. From left to right, it includes a search bar for concept lookup, tabs for navigating between ontology instances, and a plus button to open the ontology selection panel (Figure 4).
The first header element is the search bar. Users can look for concepts by their labels and directly jump to them. Users’ input text is matched against concepts labels in the current system. Expert users already familiar with an ontology can quickly navigate to known concepts using the search bar. As well, users can check whether a concept exists in a loaded ontology. For instance, in a biomedical ontology, users can search for Heart to verify its presence. If no results are returned, the concept is likely absent. Related results, such as Heart block, are also shown, helping users discover similar concepts. The search bar is not FuzzyVis’s primary means of navigating the ontology and is instead intended for quick repositioning to known landmark concepts.
The second header element consists of tabs to switch between multiple loaded ontologies. Each loaded ontology retains its own settings, highlight rules, pinned concepts, queries, and visualizations. Users can explore multiple related ontologies in parallel, or explore the same ontology using different embeddings and fuzzy operators. The tabs allow users to organize their queries into separate instances, create dedicated highlighting rules for different ontology regions, and compare the effectiveness of different embeddings. For large ontologies, having multiple instances of an ontology helps users explore different regions without having to remember exactly where they were or repeatedly update the concept collection. Users add new tabs with the ontology selection menu.
The last header element opens the ontology selection menu (Figure 4. Users can choose from preprocessed ontologies or upload their own. Uploaded files are sent to the back-end for processing. The back-end stores each ontology and its embedding in separate temporary databases. Preprocessed ontologies each reuse their own static database. The embedding options act the same as the ontology ones. Users can upload different embeddings to evaluate their effectiveness or use the pre-made ones. The last options are for configuring the fuzzy logic operators used for query resolution. Depending on the fuzzy DL reasoner used aligning an embedding with the matching fuzzy operators may be necessary. The options of product, Gödel, or Łukasiewicz t-norms cover the majority fuzzy reasoners and even when mismatched with the reasoner either of the three will often work.

### v3

@@ Abstract
The visual interface supports browsing, query composition, and partial search without requiring formal syntax. By combining fuzzy semantics with embedding-based reasoning, FuzzyVis enables flexible interpretation, efficient computation, and exploratory learning. A usage scenario demonstrates how FuzzyVis supports subtle information needs and helps users uncover relevant concepts in large, complex ontologies.
@@ 1 Introduction
To implement our querying approach, we represent ontologies as fuzzy ontologies with fuzzy interpretations, where each concept is treated as a fuzzy set assigning a membership degree to each element in the interpretation domain. Given an ontology, we use existing fuzzy reasoners to construct such interpretations that approximately satisfy the ontology’s axioms. Based on these interpretations, we introduce a novel method for generating ontology embeddings, where each concept is represented as a vector of membership degrees over a fixed, ordered set of domain elements. This allows us to precompute embeddings for all primitive concepts offline. At query time, when a user defines a composite concept using conjunction, disjunction, or negation over known primitives, its embedding is computed by applying the corresponding fuzzy operations element-wise to the membership vectors. This compositional capability enables flexible query construction and is not supported by existing ontology embedding methods. The resulting embeddings also support efficient similarity computations (e.g., cosine similarity) between user-defined concepts and existing primitive concepts in the ontology, enabling approximate matching even when no exact concept exists.
@@ 3.1 Ontologies: Syntax and Semantics
In our setting, we are more concerned with querying ontologies to retrieve individuals or concepts that match a user’s information needs. Common query types include conjunctive queries (CQs), unions of conjunctive queries (UCQs), and recursive extensions that support reachability-like conditions. DL query languages such as SPARQL-DL [55], nRQL [32], and EQL-Lite [16] combine logical reasoning with structured query syntax. However, these languages require familiarity with formal syntax and detailed knowledge of the ontology’s structure, and their evaluation is often computationally expensive. This limits their suitability for users seeking intuitive, flexible, and interactive exploration.
@@ 4 Ontology Querying using Fuzzy Ontology Embeddings
Essentially, each domain element acts as a point in the ontology’s semantic space. For each concept , the index corresponding to is a score of how ’close’ or ’inside’ an element is to . The concept embeddings are vectors of these scores, positioning each concept in the semantic space relative to the domain elements. This allows us to represent each concept not as arbitrary points, but as regions capturing a semantic profile. Given a set of domain elements representing a hierarchical ontology, a root concept would have high membership scores to all the domain elements. The embedding of this root would reflect that the concept contains everything in the ontology’s semantic space. Whereas, a leaf concept would likely only have sparse membership to a few or a single domain element and the leaf’s embedding would represent this.
@@
Scalable and optimized retrieval: Since all concept representations are vectors, similarity search over primitive concepts can be delegated to optimized vector databases (e.g., FAISS [39], Qdrant, or Chroma [20]). These systems support fast top- retrieval and efficient indexing, enabling responsive and semantics-aware retrieval for exploratory query interfaces.
4.1 -Embeddings for Hierarchical Ontologies
In many practical domains such as biomedicine, ontologies take the form of hierarchies or trees, where each concept has a parent concept (possibly more), some number of children, and there are no cycles (e.g., HPO [31], PO [38]). For such ontologies, we can construct a synthetic but semantically meaningful fuzzy interpretation that captures hierarchical relationships and allows us to generate expressive embeddings efficiently. We refer to these synthetic embeddings as -embeddings.
@@
This method provides an effective and interpretable way to construct embeddings from hierarchical ontologies, capturing both local concept similarity and the global structure. It also allows control over smoothness and generalization via the decay parameter , making it a practical tool for building fuzzy ontology embeddings in common tree-structured settings.
@@ 5.1 FuzzyVis Architecture and Implementation
The system workflow begins with the user uploading ontology data in standard formats such as .owl or .ttl. The back-end then processes and stores the ontology, making it available for the front-end. Once loaded in the front-end, users can explore the ontology through interactive views and visualizations. This exploration is supported by panels and controls that allow users to manipulate views, adjust parameters, and construct queries. As users interact with the system, FuzzyVis’s front-end dynamically retrieves relevant ontology data from the back-end.
While the front-end handles user interaction and visualization, all computations—including query evaluation and similarity search—are performed on the back-end. Embeddings for primitive concepts are stored in a Chroma vector database, as described in Section 4. This enables fast query resolution, especially for ontologies with thousands of concepts and/or large embeddings.
@@ 5.2 Back-End Server
The back-end consists of four main components: the ontology loader, ontology database (SQLite), vector (embedding) database, and the query answering component. The ontology loader handles uploaded ontologies—pre-uploaded or submitted via the front-end. It parses ontology files and stores their contents in an ontology database using the OwlReady2 python library [42]. During this process, additional metadata and statistics—such as subtree sizes and number of children—are computed and stored. The vector database stores the embeddings for each primitive concept which are generated using either a fuzzy ontology reasoner or our -embeddings as we explained in Section 4.1. Just like for ontologies, users can submit embedding vectors or use pre-uploaded ones. The vector database is separate from the ontology database to perform quick similarity calculation and top- retrieval. Our implementation uses Chroma for its vector database. The query answering component computes embeddings for user-defined composite concepts using the embeddings of primitive concepts from the vector database. The vector database uses these composite concept embeddings to retrieve similar primitive concepts. A FLASK server API handles communication between the back- and front-end.
@@ 5.3 Front-End Interface
The front-end is composed of several regions consisting of panels and controls that support interactive ontology exploration. Figure 2 shows an overview of the front-end interface as seen by the user. The primary visualization takes up the central region of the interface and is the user’s primary means of navigating the ontology (Figure 2.A). Users can change between different visualizations—treemps, nested lists, and network graphs—in this central region. The second region (Figure 2.B), the top header, provides quick access for concept search, switching between loaded ontologies, and loading ontologies. Users can manipulate the behavior and control the presentation using the visualization controls in the leftmost region (Figure 2.C). The first of the rightmost regions (Figure 2.D) is the concept panels where users can store concepts for future use and look at concept details. The second rightmost region (Figure 2.E) is the query building panels where users define queries and examine their results. FuzzyVis ’s front-end is compatible with all major web browsers.
@@ 5.3.1 Primary Visualizations
FuzzyVis ’s central region contains the primary visualization (Figure 2.A). Here, users see a representation of the ontology to aid them in exploration and understanding. Each component of FuzzyVis serves to modify, control, and navigate the primary visualization, with the back-end providing any information needed by the primary visualization. The primary visualization can take the form of any suitable representation of the ontology. Visualization should meet the following criteria:
@@
Classical visualizations such as nested lists, icicle charts, network graphs, and treemaps only partially meet these requirements. In fact, no single static visualization can satisfy every requirement. Network graphs become unreadable with large numbers of concepts, nested lists waste space, and treemaps struggle to convey secondary information. Our approach addresses these limitations by combining the primary visualization with secondary panels for additional details and utilizing dynamic interactive visualizations. To illustrate how FuzzyVis meets these requirements, we describe the nested-treemap version of the primary visualization.
Nested-treemaps represent concepts as rectangles, place sibling concepts adjacent to each other, and nest child concepts inside their parent’s area to indicate hierarchy. They are well-suited for ontology visualization because they clearly distinguish concepts, are consistent, space-efficient, and structure-revealing. Extending nested-treemaps to be dynamic and interactive allows for the display of secondary properties and addresses the problem of excessive concept density. Figure 3 shows FuzzyVis ’s nested-treemap (hereafter, the treemap), which displays concepts within an adjustable distance of a selected concept. As users navigate, the treemap updates to show related concepts and integrates with other FuzzyVis front-end components to support exploration.
@@
Selecting a concept in the treemap refocuses the visualization onto it, allowing users to drill down into its subtree or navigate upward to parent concepts. Users can also drag concepts to other components and pin them for later reference. Pinning creates visual marks that aid in orientation by acting as signposts. These marks help identify when a concept appears in multiple subtrees. For example, in HPO the Cardiac valve calcification concept is a child of both Cardiovascular calcification and Abnormal heart morphology; when pinned, it becomes more noticeable in both locations, indicating structural overlap. Without such aids, users would need to manually read labels to detect these cases. Once users have collected concepts of interest, they can create queries using the Query Building Panels to guide further exploration.
Users may want to focus on a concept while retaining its context within the ontology. Simply increasing depth causes two issues. One, deeper child concepts receive minimal space, forcing users to mouse over for tooltips to identify them. Two, densely populated subtrees create visual clutter that overwhelms users. Selecting the concept alone does not solve this, as it refocuses the visualization and removes the surrounding context.
@@
To address this, we implemented a toggle for a focus mode that enables on-the-fly visual distortions. During focus mode, when users mouse over a concept, FuzzyVis applies a discrete Cartesian distortion (often referred to as a fisheye distortion or lens), enlarging its area while preserving relative position. Within the expanded area, the subtree extends, showing the focused concept’s children; users can drill down further by mousing over these until no more space is available. If users want to explore multiple subtrees simultaneously they can lock a concept as a locus and the distortion will maintain multiple focal points. Moreover, locked loci can be transformed into a different visualization. For example, a concept cell can be converted to a node network graph or a panel showing class details. In essence, this technique directly addresses the problem of maintaining context while refocusing.
@@
Figure 3 shows an example of focus mode in use. With depth set to two and Abnormal nervous system physiology selected, no concepts other than its children would be visible. Yet, with focus mode turned on, hovering over Abnormal central motor function fills the subtree with child concepts. Afterwards, when the user mouses over the child Abnormality of coordination, the subtree is further extended to show this concept’s children (e.g., Slurred speech). This can continue until a leaf node is reached or space runs out, but the user decided this depth was sufficient and locked Abnormality of coordination as a locus. Focus mode thus offers a dynamic and space-efficient way to explore deeper while maintaining orientation. Furthermore, dynamic distortions can be applied to other kinds of visualizations, such as network graphs and charts.
Users adjust the primary visualization through the Visualization Controls (Section 5.3.3). For nested-treemaps, users can adjust depth, the label content (e.g., depth numbers), tiling methods, concept visibility, and highlighting. Such controls are essential, as different ontologies are best viewed with different settings. For instance, in HPO, Abnormality of the musculoskeletal system dominates the space when scaling is proportional to subtree size, making the smaller Abnormality of the voice nearly invisible. Equal scaling makes both visible, reducing the chance of missing concepts. Furthermore, users can toggle such options to see how the visualization changes—turning equal scaling on and off shows users which concepts take up the majority of an ontology while quickly returning to a view suitable for navigation. These adjustments enable richer exploration without restricting users to suboptimal visualizations.
@@
FuzzyVis ’s primary visualization, in conjunction with other front-end panels, provides a spatial overview of ontology structure and semantics. By supporting dynamic, interactive navigation, it helps users quickly form mental models and identify key regions of interest—an essential capability for working with complex knowledge domains where manual inspection would be slow and overwhelming. Ultimately, the primary visualization serves as the main interface through which users interact with and explore ontologies.
@@ 5.3.2 Header Bar
FuzzyVis’s header bar (Figure 2.B) contains three interactive elements: a search bar for concept lookup, tabs for switching between ontology instances, and a plus button to open the ontology selection panel (Figure 4). These allow users to load ontologies, switch among them, and perform quick keyword searches.
The search bar allows users to find concepts by label and navigate directly to them. The system matches the input text against concept labels in the current ontology, enabling rapid navigation for users familiar with the domain and quick verification of concept existence. For example, in a biomedical ontology, searching for Heart confirms whether it is present and suggests related results (e.g., Heart block) for discovery. While not intended as the primary navigation method, it is useful for repositioning to known landmark concepts.
The next header element—tabs—allows users to switch between multiple loaded ontologies. Each retains its own settings, highlight rules, pinned concepts, queries, and visualizations. Tabs support parallel exploration of related ontologies, comparison of different embeddings and fuzzy operators, or exploration of separate regions within a large ontology without losing placement. Users can organize queries, create dedicated highlight rules, and compare configurations across instances. New tabs are added via the ontology selection menu.
The plus button opens the ontology selection panel (Figure 4), where users can choose from preprocessed ontologies or upload their own. Uploaded files are processed by the back-end, which stores each ontology and its embedding in separate temporary databases; preprocessed ontologies use static databases. Embeddings can be uploaded or selected from pre-made options to evaluate effectiveness. Users can also configure fuzzy logic operators for query resolution. Depending on the fuzzy DL reasoner used, aligning an embedding with the matching fuzzy operators may be necessary. The provided product, Gödel, and Łukasiewicz t-norms cover most cases—often functioning well even when not perfectly matched to the reasoner.